OpenStudy (anonymous):

Solve: -4(2x - 1) = -6(x + 2) - 2

OpenStudy (anonymous):

-4(2x-1)=-6(x+2)-2 -8x+4=-6x-12-2 -8x+4=-6x-14 18=2x x=9

OpenStudy (anonymous):

-4(2x-1)=-6(x+2)-2 -8x+4=-6x-12-2 -8x+4=-6x-14 -8x+6x=-14-4 -2x=-18 x=-18/-2 x=9

OpenStudy (anonymous):

-9 or 9?

OpenStudy (anonymous):

9